# Cover times with stochastic resetting

### v1

@@
Cover times quantify the speed of exhaustive search. In this work, we compute exactly the mean cover time associated with a one-dimensional Brownian search under exponentially distributed resetting. We also approximate the moments of cover times of a wide range of stochastic search processes in $d$-dimensional continuous space and on an arbitrary discrete network under frequent stochastic resetting. These results hold for a large class of resetting time distributions and search processes including diffusion and Markov jump processes.
SL: Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (slinn.math@gmail.com). SL was supported by the National Science Foundation (Grant No. 2139322)
SDL: Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (lawley@math.utah.edu). SDL was supported by the National Science Foundation (CAREER DMS-1944574 and DMS-2325258).
@@ 1. Introduction
An exhaustive search is one in which an entire state space is explored by a searcher. Such search processes are ubiquitous in technology and the natural world, ranging from computer search algorithms to animal foraging or pathogen hunting. In some cases the state space of interest, or target region, may be a proper subset of the searcher domain but it may also be the entire domain. Whatever the target region, we call the time taken to complete such a search its cover time [1].
To be precise, denote the position of a stochastic process in a domain $M\subseteq\mathbb{R}^{d}$ by $X=\{X(t)\}_{t\geq 0}$. If this stochastic ‘searcher’ has detection radius $R>0$, then the region it explores by time $t\geq 0$ is
$S_{0}(t):=\cup_{s=0}^{t}B(X(s),R)\subseteq M,$
@@
$T_{0}:=\inf\{t>0:U_{T}\subseteq S_{0}(t)\}.$
The bulk of previous work on cover times concerns the case that the searcher detection radius vanishes compared to the size of the domain. In continuous space, it has been shown that cover times of Brownian motion diverge as the detection radius goes to zero [2]:
@@
Throughout, $f\sim g$ denotes $f/g\to 1$. In discrete space, it was shown that cover times of random walks follow a Gumbel distribution and diverge in the limit of a large target region [3].
More recently, others have considered cover times of multiple simultaneous searchers of fixed detection radius in the many searcher limit. In continuous space, the $m$th moment of the cover time of $U_{T}\subseteq M$ for $N\gg 1$ Brownian searchers, $T(N)\geq 0$, satisfies
@@
where $D>0$ denotes searcher diffusivity and $L>0$ denotes the distance the searcher must travel to reach the furthest part of the target region [4]. Moreover, in discrete space, these moments are given by
$\mathbb{E}((T(N))^{m})\sim\frac{\kappa_{m}}{N^{m/d}}\quad\text{as}N\to\infty,$
where $d\geq 1$ is the fewest steps the searcher must take to reach the furthest part of the target region and $\kappa_{m}$ depends on the jump rates along this path [5]. An important observation in the $N\gg 1$ searcher case is that the cover times can be closely approximated with merely some information about the geodesic paths from the initial position of the searcher to the furthest part(s) of the target region. The same can be said of the work herein.
@@
Not only is stochastic resetting a useful formalism to model physical phenomena [7, 8, 9], it can also expedite certain search processes [10, 11, 12]. For instance, consider the mean escape time of a Brownian searcher with diffusivity $D>0$ on the half line. It is well-known that without stochastic resetting the mean first passage time to the origin is infinite. That is, denoting its position at time $t\geq 0$ by $\{X(t)\}_{t\geq 0}$ and supposing $X(0)=x_{0}>0$, we have that
$\mathbb{E}[\text{inf}\{t>0:X(t)\leq 0\}]=\infty.$
@@
However, if we now impose that this searcher instantaneously resets to its initial position according to an exponential random variable with rate $r>0$, the mean first passage time to the origin is finite,
$\mathbb{E}[\text{inf}\{t>0:X(t)\leq 0\}]=\frac{1}{r}(e^{x_{0}\sqrt{r/D}}-1)<\infty,$
and attains a minimum for a strictly positive resetting rate [13].
@@
Similar to mean first passage times, we show that stochastic resetting can be used to reduce MCTs in continuous space. Moreover, with a well-defined notion of resetting rate, $r>0$, we show that the frequent resetting asymptotics of the moments of cover times in discrete and continuous space depend only on the geodesic path(s) from the initial (resetting) position of the searcher to the furthest part(s) of the target region. Our results hold for any fixed detection radius. While these results are exact only in the frequent resetting limit, we can compute exactly the mean cover time with any resetting rate $r>0$ in the special case of a one-dimensional Brownian searcher.
@@
All these results are made precise in the rest of the paper, which is organized as follows. In section 2, we consider the one-dimensional Brownian cover time problem with resetting. In section 3, we consider the $d$-dimensional continuous space cover time problem with resetting. In section 4, we consider the cover time problem with resetting on an arbitrary discrete network. We conclude by discussing the scope of the results and proposing directions of future work.
2. Exact mean cover times for a resetting Brownian searcher in $\mathbb{R}$
Consider a Brownian searcher with diffusivity $D>0$ who resets according to exponential random variable with rate $r>0$ to its initial position. We denote its position at time $t\geq 0$ by $X_{r}(t)$ and denote its first hitting time to a set $\Omega$ with initial position $x_{0}$ and resetting position $x_{r}$ by $\tau_{r}(\Omega|x_{0},x_{r})$. If $x_{0}=x_{r}=0$, we abbreviate $\tau_{r}(\Omega)\equiv\tau_{r}(\Omega|0,0)$.
In this section, we determine the mean cover time of the interval $[-a,b]$ with $a>0$ and $b>0$. We consider two cases: In the first case the searcher is unconstrained on the real line and in the second case the searcher is confined to the interval $[-a,b]$ with reflecting boundary conditions. In section 3, we ascribe a detection radius to the searcher and study this system in higher dimensions and more general domains.
2.1. Unconstrained search
Consider first that the searcher is unconstrained on $\mathbb{R}$. It follows from the strong Markov property that the cover time of the interval $[-a,b]$ can be described by the sum of the exit time from $(-a,b)$ and the exit time from the half line with initial position $X_{r}(0)=a+b$ and resetting position either $a>0$ or $b>0$ depending on which interval boundary was first contacted. The mean interval exit time is known [14],
(3) $\epsilon_{a}$ $:=\mathbb{P}(\tau_{r}(\{-a\})<\tau_{r}(\{b\}))=\frac{\sinh(b\alpha)}{\sinh(a\alpha)+\sinh(b\alpha)},$
(4) $\epsilon_{b}$ $:=1-\mathbb{P}(\tau_{r}(\{-a\})<\tau_{r}(\{b\}))=\frac{\sinh(a\alpha)}{\sinh(a\alpha)+\sinh(b\alpha)}.$
Denote the cover time by $T_{r}>0$. Combining equations 1, 3, 4 and 2 yields the MCT of the interval $[-a,b]$ by a resetting Brownian searcher,
(5) $\begin{split}\mathbb{E}(T_{r})&=\mathbb{E}(\tau_{r}(\{-a,b\}))+\epsilon_{a}\mathbb{E}(\tau_{r}(\{0\}|a+b,b))+\epsilon_{b}\mathbb{E}(\tau_{r}(\{0\}|a+b,a))\\ &=\frac{e^{(a+b)\alpha}+e^{2a\alpha}+e^{2b\alpha}-e^{a\alpha}-e^{b\alpha}-1}{r\left(e^{a\alpha}+e^{b\alpha}\right)}.\end{split}$
Notice in Figure 1 that for finite resetting rate $r$ the mean cover time is finite. This is in stark contrast to the analogous cover time problem without resetting; it is well-known that the mean first passage time to a point not equal to $X(0)$ is infinite and, hence, so too is the mean cover time of an interval of nonzero length. Observe further that the MCT of the interval $[-a,b]$ by a resetting Brownian motion on the (unconstrained) real line achieves a unique minimum with a positive resetting rate. This ‘optimal’ resetting rate, which is found by computing the unique zero of the $r$-derivative of (5), does not have an explicit analytical solution, but notice in Figure 1 how closely the frequent resetting result approximates the exact result.
Taking $r\to\infty$ in (5) reveals that the MCT depends on the relative sizes of $a>0$ and $b>0$. For instance, if $a<b$, one can verify that the large $r$ behavior of (5) behaves like
(6) $\mathbb{E}(T_{r})\sim\frac{1}{rp}\quad\text{as }r\to\infty$
where $p:=\text{exp}(-b\alpha)$ is the probability of hitting the further boundary point. In this case it is straightforward to compute the minimum MCT of the asymptotic result in (6) of $b^{2}e^{2}/4D$ at $r=4D/b^{2}$, which converges to the exact result exponentially fast in the square root of the resetting rate. Hence, the asymptotic result in (6) closely approximates the true optimum even for modest resetting rates. We find similar large $r$ behavior in much more general settings in section 3.
We briefly address the MCT of a resetting Brownian motion on a symmetric interval, i.e. the case $|a|=b$. Repeating the analysis above yields the same asymptotic result as in (6) with a prefactor of $3/2$. The same phenomenon can be observed in the case of a constrained search, which we consider in subsection 2.2 below. We give an intuitive explanation for this prefactor in the discussion.
2.2. Constrained search
Consider the same process but on the closed interval $[-a,b]$ with reflecting boundary conditions. The cover time of the interval can now be described by the sum of the hitting time of $x\in\{-a,b\}$ and the hitting time of $x=0$ on the closed interval $[0,a+b]$ with initial position $X_{r}(0)=a+b$ and resetting position either $a>0$ or $b>0$ depending on which interval boundary was first hit. The mean of the first term was computed in the previous section. Below we determine the mean time of the second term.
$Q_{r}(x,t):=\mathbb{P}(\tau_{r}(\{-a\})>t|X_{r}(0)=x).$
Figure 1. MCTs of the interval $[-1,2]$ for a resetting Brownian motion with diffusivity $D=1$ and $L=2$ denoting the furthest target point from the searcher initial and resetting position. Resetting is exponentially distributed with rate $r>0$. See sections 2 and 3 for details.
with boundary conditions $Q_{0}(-a,t)=0$ and $\frac{\partial Q_{0}(x,t)}{\partial x}\big{|}_{x=b}=0$. Denote the Laplace transform in time of $Q_{r}(x,t)$ by $q_{r}(x,s)$. The Laplace transform of (7) satisfies the ordinary differential equation
$sq_{0}(x,s)-1=Dq^{\prime\prime}_{0}(x,s)$
with boundary conditions $q_{0}(-a,s)=0$ and $\frac{\partial q_{0}(x,s)}{\partial x}\big{|}_{x=b}=0$. We substitute the solution of this boundary value problem into a known expression of $q_{r}(x,s)$ in terms of $q_{0}(x,s)$ (which can be derived by way of a renewal equation, see [13]),
(8) $q_{r}(x_{0},s)=\frac{q_{0}(x_{0},r+s)}{1-rq_{0}(x_{r},r+s)},$
and finally evaluate (8) at $s=0$ with $x_{0}=b$ and $x_{r}=0$ to find that
Combining equations 9 and 10 with the mean time of the first term and the relevant hitting probabilities computed in the previous system yields the MCT of the closed interval $[-a,b]$ by a resetting Brownian searcher,
(11) $\begin{split}\mathbb{E}(T_{r})=\frac{1}{r}&\left[\frac{\sinh((a+b)\alpha)}{\sinh(a\alpha)+\sinh(b\alpha)}-1\right]\dots\\ &\qquad+\frac{\epsilon_{a}}{r}\left[\frac{(e^{(a+b)\alpha}-1)^{2}}{e^{b\alpha}(e^{2a\alpha}+1)}\right]+\frac{\epsilon_{b}}{r}\left[\frac{(e^{(a+b)\alpha}-1)^{2}}{e^{a\alpha}(e^{2b\alpha}+1)}\right].\end{split}$
$\mathbb{E}(T_{r})\sim\frac{1}{rp}\quad\text{as }r\to\infty$
where $p:=\text{exp}(-b\alpha)$ is again the probability of hitting the further boundary point before a resetting event. We illustrate this behavior in Figure 1. Similar to the unconstrained case, the asymptotic result here closely approximates the true optimum even for modest resetting rates.
@@ 3.1. Definitions and main result
Now consider a stochastic searcher in the domain $M\subseteq\mathbb{R}^{d}$ with fixed detection radius $R>0$ which randomly resets to its initial position. We are most interested in the case of Brownian search, which we later show satisfies the minimal assumptions we impose on the search process. Denoting searcher position at time $t\geq 0$ by $X_{r}(t)$, the region detected by the searcher at time $t\geq 0$ given a metric $L:M\times M\to[0,\infty)$ is the closed ball of radius $R>0$ centered at $X_{r}(t)\in M$,
@@
For Brownian search, $L$ is merely the standard Euclidean distance. We assume the initial and resetting position of the searcher is $X_{r}(0)=0$ and denote by $U_{0}$ the region it initially detects,
$U_{0}:=B(X_{r}(0),R)=B(0,R).$
We consider resetting processes, $\sigma>0$, that belong to the following family of random variables: Let $Y>0$ be a strictly positive random variable that does not depend on $r>0$ and has unit mean,
@@
(12) $\sigma:=Y/r.$
where we call $r>0$ the resetting rate. If $Y$ is a unit rate exponential random variable, then $\sigma$ is exponentially distributed with rate $r>0$.
The region covered by the resetting searcher up through time $t>0$ is
@@
$T_{r}:=\inf\{t>0:U_{T}\subseteq S_{r}(t)\}.$
We denote by $U_{T}^{\text{far}}\subseteq U_{T}$ the set of furthest target points, which means that
$L(x,X_{r}(0))>L(y,X_{r}(0))\quad\text{for all }x\in U_{T}^{\text{far}}\text{ and }y\in U_{T}\backslash U_{T}^{\text{far}}.$
Analogous to section 2 we define the first hitting time of $X_{r}(t)$ to a set $\Omega\subseteq M$,
$\tau_{r}(\Omega):=\inf\{t\geq 0:X_{r}(t)\in\Omega\}.$
We remark that throughout this work $T$ denotes a cover time and $\tau$ denotes a hitting time. If $\Omega\equiv B(x,R)$, then $\tau_{r}(\Omega)$ is equivalent to
$\tau_{r}(B(x,R))=\inf\{t>0:X_{r}(t)\in B(x,R)\}=\inf\{t>0:x\in S_{r}(t)\}.$
Finally, denote by $p_{\Omega}$ the probability of a stochastic searcher in $M\subseteq\mathbb{R}^{d}$ with initial position $X_{0}(0)=0$ hitting the set $\Omega\subseteq M$ before a resetting event at time $\sigma>0$,
$p_{\Omega}:=\mathbb{P}(\tau_{0}(\Omega)<\sigma).$
A key assumption in the results of this section is that for two sets $\Omega_{0},\Omega_{1}\subset M$,
@@
where the distance between two sets $\Omega_{0},\Omega_{1}\subset M$ is given by
$L(\Omega_{0},\Omega_{1}):=\inf_{x\in\Omega_{0},y\in\Omega_{1}}L(x,y).$
@@
$\ln p_{\Omega}\sim-\sqrt{L(\Omega_{0},X_{r}(0))^{2}r/D}\quad\text{as }r\to\infty$
where $D>0$ is the diffusivity (see Theorem 2 in [15]).
@@
Under the assumptions of section 3, let $p^{*}=p^{*}(r)$ be any function of $r>0$ satisfying $p^{*}\sim p_{B(x^{*},R)}$ as $r\to\infty$ for any $x^{*}\in U_{T}^{\text{far}}$. Then
$\liminf_{r\to\infty}\mathbb{E}[(rp^{*}T_{r})^{m}]\geq m!\quad\text{as }r\to\infty$
for integers $m\geq 1$. Further,
@@
In words, Theorem 1 says that in the frequent resetting limit we can bound a certain scaling of the cover time that depends only on the resetting rate and data on the set of furthest target points.
3.2. Proofs
Before we prove Theorem 1, we introduce a useful lemma which was recently proven (see Theorem 1 in [15]):

### v2

@@
Cover times quantify the speed of exhaustive search. In this work, we approximate the moments of cover times of a wide range of stochastic search processes in $d$-dimensional continuous space and on an arbitrary discrete network under frequent stochastic resetting. These approximations apply to a large class of resetting time distributions and search processes including diffusion, run-and-tumble particles, and Markov jump processes. We illustrate these results in several examples; in the case of diffusive search we show that the errors of our approximations vanish exponentially fast. Finally we derive a criterion for when endowing a discrete state search process with minimal stochastic resetting reduces the mean cover time.
SL (corresponding author): Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (slinn.math@gmail.com).
SDL: Department of Mathematics, University of Utah, Salt Lake City, UT 84112 USA. (lawley@math.utah.edu).
Cover times measure the time taken by a stochastic searcher to find multiple targets. Cover times thus naturally extend first passage times (FPTs), which describe the time to find a single target. Recent work has shown that stochastic resetting, wherein a searcher randomly resets in space, can reduce mean FPTs. In this work, we study cover times with stochastic resetting. In particular, we approximate statistics of cover times under frequent stochastic resetting. Using recent results on FPTs, we find that such cover times resemble FPTs to the most distant targets. Quantitatively, these cover times are approximately exponentially distributed with a rate depending on the resetting rate and the distant target positions. We illustrate our results through numerous examples in $d$-dimensional continuous space and on a discrete network, which reveal that stochastic resetting can reduce mean cover times.
@@ 1. Introduction
An exhaustive search is one in which an entire state space is explored by a searcher. Such search processes are ubiquitous in technology and the natural world, ranging from computer search algorithms [1] to animal foraging [2, 3] or pathogen hunting [4]. In some cases the state space of interest, or target region, may be a proper subset of the searcher domain but it may also be the entire domain. Whatever the target region, we call the time taken to complete such a search its cover time.
To be precise, denote the position of a stochastic process in a domain $M\subseteq\mathbb{R}^{d}$ by $X=\{X(t)\}_{t\geq 0}$. If this stochastic ‘searcher’ has detection radius $R>0$, then the region it explores by time $t>0$ is
$S_{0}(t):=\cup_{s=0}^{t}B(X(s),R)\subseteq M$
@@
$T_{0}:=\inf\{t\geq 0:U_{T}\subseteq S_{0}(t)\}.$
The bulk of previous work on cover times concerns stochastic search processes on discrete state spaces [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. A seminal result on this topic is that, in the limit of a large target region, cover times of random walks follow a Gumbel distribution [8]. There has also been notable progress characterizing cover times on continuous state spaces by making use of various asymptotic limits. For instance, it has been shown that cover times of Brownian motion diverge as the detection radius goes to zero [16],
@@
Throughout, $f\sim g$ denotes $f/g\to 1$.
In more recent work, others have considered cover times of multiple simultaneous searchers of fixed detection radius in the many searcher limit. In continuous space, the $m$th moment of the cover time of $U_{T}\subseteq M$ for $N\gg 1$ Brownian searchers, $T(N)\geq 0$, satisfies
@@
where $D>0$ denotes searcher diffusivity and $L>0$ denotes the shortest distance the searcher must travel to reach the furthest part of the target region [17]. Moreover, in discrete space, these moments are given by
$\mathbb{E}((T(N))^{m})\sim\frac{\kappa_{m}}{N^{m/h^{*}}}\quad\text{as }N\to\infty,$
where $h^{*}\geq 1$ is the fewest steps the searcher must take to reach the furthest part of the target region and $\kappa_{m}$ depends on the jump rates along this path [18]. An important observation in the $N\gg 1$ searcher case is that the cover times can be closely approximated with merely some information about the geodesic paths from the initial position of the searcher to the furthest part(s) of the target region. The same can be said of the work herein.
@@
Not only is stochastic resetting a useful formalism to model physical phenomena [20, 21, 22], it can also expedite certain search processes [23, 24, 25, 26]. For instance, consider a Brownian searcher with diffusivity $D>0$ on the half line. It is well-known that without stochastic resetting the mean first passage time (MFPT) to the origin is infinite.
Figure 1. Representative illustration of a spatial domain $M$ (pictured, a 2D torus) with a stochastic searcher that starts at $U_{0}$. Dark purple indicates the searcher path and light purple indicates the corresponding region detected. Orange denotes the target region ($U_{T}$) and yellow denotes the shortest path (length $L$) that the searcher must travel to cover the farthest part of the target.
@@
However, if this searcher instantaneously resets to its initial position according to an exponential random variable with rate $r>0$, then the MFPT to the origin is finite. That is, denoting its position at time $t>0$ by $X=\{X(t)\}_{t\geq 0}$ and supposing $X(0)=x_{0}>0$, we have that
$\mathbb{E}[\text{inf}\{t\geq 0:X(t)\leq 0\}]=\frac{1}{r}(e^{x_{0}\sqrt{r/D}}-1)<\infty,$
which attains a minimum for a strictly positive resetting rate [27].
@@
Similar to MFPTs, we show that stochastic resetting can be used to reduce MCTs, though the resetting rates that achieve these reductions may differ between the two. Moreover, we show that the frequent resetting asymptotics of the moments of cover times in discrete and continuous space depend only on the geodesic path(s) from the initial (resetting) position of the searcher to the furthest part(s) of the target region. Our results hold for any fixed detection radius. While these results are exact only in the frequent resetting limit, we compute exactly the MCT with any resetting rate $r>0$ in the cases of a one-dimensional Brownian searcher and run-and-tumble particle.
@@
These results are made precise in the rest of the paper, which is organized as follows. In section 2, we consider the $d$-dimensional continuous space cover time problem with resetting. In section 3, we consider the cover time problem with resetting on an arbitrary discrete network. Section 4 illustrates these general results in several examples. We conclude by discussing the scope of the results and proposing directions of future work.
@@ 2.1. Definitions and main result
Now consider a stochastic searcher in the domain $M\subseteq\mathbb{R}^{d}$ equipped with a metric $L:M\times M\to[0,\infty)$ and with fixed detection radius $R>0$ which randomly resets to its initial position. The searcher may be Brownian, though other stochastic processes also satisfy the minimal assumptions imposed on the search process (see section 4). Denoting the searcher position at time $t\geq 0$ by $X_{r}(t)$, the region detected by the searcher at time $t\geq 0$ is the closed ball of radius $R>0$ centered at $X_{r}(t)$,
@@
The subscript $r$ of $X_{r}$ denotes the resetting rate, which we formalize below. Denote by $U_{0}$ the region the searcher initially detects,
$U_{0}:=B(X_{r}(0),R).$
We consider resetting processes, $\sigma>0$, that belong to the following family of random variables: Let $Y>0$ be a strictly positive random variable with unit mean,
@@
(1) $\sigma:=Y/r$
where we call $r>0$ the resetting rate. If $Y$ is a unit rate exponential random variable, then $\sigma$ is exponentially distributed with rate $r$.
The region covered by the resetting searcher up through time $t\geq 0$ is
@@
$T_{r}:=\inf\{t\geq 0:U_{T}\subseteq S_{r}(t)\}.$
To exclude trivial cases, we assume $U_{T}\not\subseteq U_{0}$. See Figure 1 for an illustration.
Denote the first hitting time of the resetting searcher to a set $\Omega\subseteq M$ with initial position $X_{r}(0)=x_{0}$ and resetting position $x_{r}$ by
(2) $\tau_{r}(\Omega|x_{0},x_{r}):=\inf\{t\geq 0:X_{r}(t)\in\Omega\}.$
If $x_{0}=x_{r}$, we abbreviate (2) by $\tau_{r}(\Omega)$.
Assume that for any union of balls with strictly positive radii that does not contain $U_{0}$, the search time without resetting $\tau_{0}(\Omega)$ is strictly positive. If $\Omega=B(x,R)$, then $\tau_{r}(\Omega)$ is equivalent to
$\tau_{r}(B(x,R))=\inf\{t\geq 0:X_{r}(t)\in B(x,R)\}=\inf\{t\geq 0:x\in S_{r}(t)\}.$
We remark that throughout this work $T$ denotes a cover time and $\tau$ denotes a hitting time.
Denote by $p_{\Omega}$ the probability of a stochastic searcher in $M\subseteq\mathbb{R}^{d}$ with initial position $X_{0}(0)$ hitting a union of balls with strictly positive radii $\Omega\subseteq M$ before time $\sigma>0$,
(3) $p_{\Omega}:=\mathbb{P}(\tau_{0}(\Omega)<\sigma).$
To exclude trivial cases, we assume $p_{\Omega}>0$ for all $r>0$ and $\Omega\subset M$.
A key assumption of this section is that for any two unions of balls with strictly positive radii $\Omega_{0},\Omega_{1}\subset M$, we have the following,
@@
where the distance between two sets $\Phi,\Psi\subset M$ is given by
$L(\Phi,\Psi):=\inf_{x\in\Phi,y\in\Psi}L(x,y).$
@@
(5) $\ln p_{\Omega}\sim-\sqrt{L(\Omega,X_{r}(0))^{2}r/D}\quad\text{as }r\to\infty$
where $D>0$ is the diffusivity and $L$ is the standard Euclidean distance (see Theorem 2 in [28]). See subsections 4.3 and 4.4 for specific examples of this behavior.
Finally, we assume that there exists a nonempty set $U_{T}^{\text{far}}\subseteq U_{T}$ such that if $x^{*}\in U_{T}^{\text{far}}$ and $y\in U_{T}$ then
$L(y,X_{r}(0))\leq L(x^{*},X_{r}(0)).$
In words, $U_{T}^{\text{far}}$ is the set of target points which are farthest from the initial searcher position.
@@
Under the assumptions subsection 2.1,
$\liminf_{r\to\infty}\mathbb{E}[(rp_{B(y,R)}T_{r})^{m}]\geq m!\quad\text{as }r\to\infty$
for any $y\in U_{T}$ and integer $m\geq 1$. Further, for any $x^{*}\in U_{T}^{\text{far}}$,
@@
Roughly speaking, Theorem 1 gives the following bounds on the moments of the cover time in the frequent resetting regime,
(6) $\frac{m!}{(rp_{B(x^{*},R)})^{m}}\lesssim\mathbb{E}[(T_{r})^{m}]\lesssim\frac{m!}{(rp_{B(x^{*},(1-\varepsilon)R)})^{m}}.$
Hence the asymptotic behavior of $T_{r}$ is determined by the success probability associated with the furthest part of the target. The approximate behavior in (6) is clear in the case of a one-dimensional Brownian motion as demonstrated by (32) in section 4.1. We note that Theorem 1 still holds if the success probabilities are replaced by asymptotically equivalent quantities.
Finally observe that through (6) we can approximate the ‘optimal’ resetting rate that minimizes the MCT. Namely, by finding the zero of the $r$-derivative of the first relation in (6) with $m=1$, we attain the following general approximation,
(7) $\frac{1}{r_{\textup{opt}}}\approx-\frac{\textup{d}}{\textup{d}r_{\textup{opt}}}\ln p_{B(x^{*},R)}(r_{\textup{opt}}).$
In section 4 we compute the approximate optimum in various examples and compare them to exact values.
2.2. Proof
The proof of Theorem 1 relies on the following lemma, which was proven in [28] (see Theorem 1 in [28]).